OpenStudy (anonymous):

3. What is the factored form of x2 – 196? (1 point) (x – 14)2 (x + 14)2 (x – 28)(4x + 7) (x – 14)(x + 14) 4. What is the factored form of 9x2 – 64? (1 point) (3x – 8)2 (3x + 8)2 (3x – 8)(3x + 8) (9x – 8)(x + 8) 5. What is the factored form of 49x2 – 56x + 16? (1 point) (7x – 4)2 (7x + 4)(7x – 4) (7x + 4)2 (7x – 8)2

OpenStudy (anonymous):

x^2 - 14^2 = (x – 14)(x + 14) 9x2 – 64 = (3x)^2 - 8^2 = (3x -8) ( 3x + 8) 49x2 – 56x + 16 = (7x)^2 - 4^2 = (7x - 4) ( 7x + 4)

OpenStudy (anonymous):

THANK UU :P